two tanks hold a toatl of 45 gallons of toxic solvent. One tank holds 6 galoons more than twice the amount in the other. how many gallons does each tank hold.
Answer by mananth(16949)

let one tank hold x gallons
the other tank hold 2x+6 gallons
..
together they hold 45 gallons
..
x+2x+6=45
3x+6=45
add-6 to both sides
3x+6-6=45-6
3x=39
x= 39/3
x=13 gallons
the other tank will hold 45-13= 32 gallons